# Mean field equation and Green’s function on the Tate curve

### 16 Mar 2026 v1

@@ 4.1 Convergence to the Tate curve
In this part, we will show that the solution on the finite quotient will converge to the solution on the Tate curve, thus we can get the existence of the solution. From the numerical experiments, we can indeed observe that the solution on the finite quotient is convergent. Next, we will prove it rigorously.
@@
When $\rho\leq\begin{cases}\frac{p-2}{p}(1-p^{-m}),\ &p>2\\ \frac{1-2^{-m}}{4},&p=2\end{cases}$, then there exists a solution $u$ satisfying (3.1).
@@
For convenience, we denote $X_{m}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ with the Haar measure $d\mu=\frac{dx}{|x|_{p}}$ and $X_{m,d}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ with the Haar measure $d\mu_{d}=p^{-d}d\delta$. Clearly, $\mu_{d}\overset{*}{\rightharpoonup}\mu$ as $d\rightarrow\infty$. Let $D_{d}$ denote the operator on the finite quotient. Consider
@@
$\int_{X_{m,d}}|D_{d}u_{d}|d\mu_{d}\leq 2\rho$ (4.3)
Define measures $\nu_{d}=e^{u_{d}}\mu_{d}$ and $\omega_{d}=D_{d}u_{d}\mu_{d}$ on $X_{m,d}$. As $X_{m,d}\hookrightarrow X_{m}$, so they can be seen as the restriction of measure $\widetilde{\nu_{d}}$ and $\widetilde{\omega_{d}}$ on $X_{m}$, i.e. $\widetilde{\nu_{d}}|_{X_{m,d}}=\nu_{d}$ and $\widetilde{\omega_{d}}|_{X_{m,d}}=\omega_{d}$.
Since $X_{m}$ is compact, we have $C(X_{m})^{*}=BV(X_{m})$, so by the Banach-Alaoglu theorem, $BV(X_{m})$ is weak-* compact. Then from (4.2) and (4.3), we have $\widetilde{\nu_{d}}||_{BV}=1$ and $\widetilde{\omega_{d}}||_{BV}\leq 2\rho$, so there exists $\nu$, $\omega\in BV(X_{m})$ and a subsequence $d_{k}$ such that
@@
as $\nu_{d}<<\mu_{d}$ and $\frac{d\nu_{d}}{d\mu_{d}}=e^{u_{d}}\leq C_{p}$ uniformly in $d$ by Lemma 3.9, we have $\nu<<\mu$, denoting $\nu=e^{\widetilde{u}}\mu$ for some function $\widetilde{u}$.
Take any locally constant function $\phi$ on $X_{m}$, so when $d$ is large enough, $\phi$ is well-defined on $X_{m,d}$ as $X_{m}$ is compact. Multiply (4.1) by $\phi$ and integrate over $X_{m,d}$, we have
$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}+\rho\int_{X_{m,d}}e^{u_{d}}\phi d\mu_{d}=\rho\phi(y)$
@@
$v_{d}(x)=\int_{X_{m,d}}G_{d}(x,y)(-\rho e^{u_{d}(y)}+\frac{\rho}{V})d\mu_{d}(y)$
Since $G(x,y)\sim\frac{p(p-1)}{p+1}\text{log}_{p}d(x,y)$ as $d(x,y)\rightarrow 0$ [14], and $X_{m}$ is compact, then
@@
and we have $\int_{B_{\epsilon}(x)}|G(x,y)|d\mu(y)\leq C\epsilon|\text{ln}\epsilon|$, so $G\in L^{1}(X_{m})$. and we notice that $\rho e^{u_{d}(y)}+\frac{\rho}{V}$ is bounded, we have
@@
so $v_{d}\in L^{\infty}(X_{m,d})$.
@@
$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}=$ $\int_{X_{m,d}}(D_{d}\phi)u_{d}d\mu_{d}$
@@
$|\mathrm{I}|=$ $\left|\int_{B_{\epsilon}(y)}(D_{d}\phi)u_{d}d\mu_{d}\right|$
$\leq$ $\rho\int_{B_{\epsilon}(y)}|D_{d}\phi(x)||G_{d}(x,y)|d\mu_{d}(x)+\int_{B_{\epsilon}(y)}|D_{d}\phi||v_{d}|d\mu_{d}$
$\leq$ $C_{1}\epsilon|\text{ln}\epsilon|+C_{2}\epsilon$
@@
$\mathrm{II}=\int_{X_{m,d}\backslash B_{\epsilon}(y)}(D_{d}\phi)u_{d}d\mu_{d}=\int_{X_{m,d}\backslash B_{\epsilon}(y)}(D\phi)u_{d}d\mu_{d}$
for sufficiently large $d$. And in $X_{m,d}\backslash B_{\epsilon}(y)$, $|u_{d}|\leq C_{\epsilon}$, so there exists a subsequence $u_{d_{k}}\rightarrow u$ such that
$\mathrm{II}\rightarrow\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)ud\mu$
letting $\epsilon\rightarrow 0$ we obtain
$\int_{X_{m,d_{k}}}(D_{d_{k}}u_{d_{k}})\phi d\mu_{d_{k}}\rightarrow\int_{X_{m}}(D\phi)ud\mu=\int_{X_{m}}(Du)\phi d\mu$
@@
What’s more,
$\int_{X_{m}}\phi d\nu=\lim_{k\to\infty}\int_{X_{m,d_{k}}}\phi e^{u_{d_{k}}}d\mu_{d_{k}}=\int_{X_{m}}\phi e^{u}d\mu$
and by definition,
$\int_{X_{m}}\phi d\nu=\int_{X_{m}}\phi e^{\widetilde{u}}d\mu$
so we have $\widetilde{u}=u$ for a.e. $x$. Then by (4.4) we get
$\int_{X_{m}}(Du)\phi d\mu+\rho\int_{X_{m}}e^{u}\phi d\mu=\rho\phi(y)$
@@
Since the solution we found converges from the solution on finite quotient, it inherits the properties of the solution on finite quotient. So the solution $u$ has those radial symmetric properties, what’s more, $u$ is monotone with respect to $d(x,y)$ at the same connected component and $u\leq C_{p}$.
When $\rho$ breaks the bound of Theorem 4.1, then there may be non-radial symmetric solutions, so that $u$ will not have the universal upper bound, this may prevent the convergence of solutions on the finite quotient, and hence the existence cannot be guaranteed.
@@ 4.2 Uniqueness of the solution
Similarly from the argument of the solution in finite quotient, we now address uniqueness on the Tate curve under the same constraints of $\rho$. Although the solution on finite quotient is unique when $\rho$ is not too large, we can’t ensure the solution on the Tate curve is unique as there may has another solution on the Tate curve but vanishes on finite quotient.
Actually, the solution is indeed unique under the same constraints of $\rho$. We will prove it by a standard method [18]. First, we show that the solution is unique when $\rho$ is very small, next we prove the linearized solution is non-degenerate under the same constraints of $\rho$, then we can get the uniqueness.
@@
$D\phi_{n}+V_{n}(x)\phi_{n}=0$
where $V_{n}(x)=\rho_{n}\frac{e^{u_{n}}-e^{v_{n}}}{u_{n}-v_{n}}$. Since $u(x)$ has a universal upper bound by Lemma 3.9, $V_{n}(x)$ converges to 0. And it is easy to see that $\phi_{n}$ uniformly converges to the eigenfunction of $D$ with the eigenvalue 0, which is the harmonic function on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$. As $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ is compact without boundary, then the harmonic function on it must be a constant function, which contradicts the fact that $\phi_{n}$ must change the sign. ∎
@@
$Df+\rho e^{u}f=0$ (4.5)
If (4.5) is nondegenerate for $\rho\leq C$, then Lemma 4.2 implies the uniqueness of the solution for such $\rho$. To address this question, we give the following Lemma.
@@
$\mu_{k}<0<\mu_{k+1}$
which means 0 is not the eigenvalue of LV(x)L-V(x). ∎
By Lemma 3.9, we can see that $0<\rho e^{u}<\rho\frac{p}{p-2}$. And from Lemma 4.3, we can see that if $\rho\frac{p}{p-2}$ is smaller than the first nonzero eigenvalue of D-D, then we can get the uniqueness, similarly, the smallest nonzero eigenvalue of D-D is
$\text{min}_{\begin{subarray}{c}1\leq k\leq m-1\\ i\geq 1\\ 0\leq j\leq m-1\end{subarray}}\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\}$ (4.6)
$=$ $\text{min}\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
@@
then we have
Theorem 4.4.
When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then the solution of (3.1) is unique.

### 08 Apr 2026 v2

@@ 4.1 Convergence to the Tate curve
When $\rho\in(0,\rho^{*})$, then there exists a solution $u$ satisfying (3.1).
@@
For convenience, we denote $X_{m}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ with the Haar measure $d\mu=\frac{dx}{|x|_{p}}$ and $X_{m,d}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ with the Haar measure $d\mu_{d}=p^{-d}d\delta$, which can be extended to the measure $\widetilde{\mu_{d}}$ on $X_{m}$. Clearly, $\widetilde{\mu_{d}}\overset{*}{\rightharpoonup}\mu$ as $d\rightarrow\infty$. Let $D_{d}$ denote the operator on the finite quotient. Consider
@@
$\int_{X_{m,d}}|D_{d}u_{d}|d\mu_{d}\leq 2\rho.$ (4.3)
Define measures $\nu_{d}=e^{u_{d}}\mu_{d}$ and $\omega_{d}=D_{d}u_{d}\mu_{d}$ on $X_{m,d}$. As $X_{m,d}\hookrightarrow X_{m}$, so they can be seen as the restriction of measure $\widetilde{\nu_{d}}$ and $\widetilde{\omega_{d}}$ on $X_{m}$, i.e. $\widetilde{\nu_{d}}|_{X_{m,d}}=\nu_{d}$ and $\widetilde{\omega_{d}}|_{X_{m,d}}=\omega_{d}$. On the other hand, the function on $X_{m,d}$ can extend to a locally constant function on $X_{m}$.
Since $X_{m}$ is compact, we have $C(X_{m})^{*}=\mathscr{M}(X_{m})$, so by the Banach-Alaoglu theorem, $\mathscr{M}(X_{m})$ is weak-* compact. Then from (4.2) and (4.3), we have $\widetilde{\nu_{d}}||_{\mathscr{M}}=1$ and $\widetilde{\omega_{d}}||_{\mathscr{M}}\leq 2\rho$, so there exists $\nu$, $\omega\in\mathscr{M}(X_{m})$ and a subsequence $d_{k}$ such that
@@
as $\nu_{d}<<\mu_{d}$ and $\frac{d\nu_{d}}{d\mu_{d}}=e^{u_{d}}\leq e^{M}$ uniformly in $d$ by Theorem 3.13, we have $\nu<<\mu$, denoting $\nu=e^{\bar{u}}\mu$ for some function $\bar{u}$.
Taking any locally constant function $\phi$ on $X_{m}$, when $d$ is large enough, $\phi$ is well-defined on $X_{m,d}$ as $X_{m}$ is compact. Multiply (4.1) by $\phi$ and integrate over $X_{m,d}$, we have
$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}+\rho\int_{X_{m,d}}e^{u_{d}}\phi d\mu_{d}=\rho\phi(y),$
and it can be rewritten as
$\int_{X_{m}}\phi d\widetilde{\omega_{d}}+\rho\int_{X_{m}}\phi d\widetilde{\mu_{d}}=\rho\phi(y).$
@@
$v_{d}(x)=\int_{X_{m,d}}G_{d}(x,y)\left(-\rho e^{u_{d}(y)}+\frac{\rho}{V}\right)d\mu_{d}(y).$
@@
And we have
$\int_{B_{\epsilon}(x)}\text{log}_{p}d(x,y)d\mu(y)=$ $\int_{x+p^{v_{p}(x)+N}\mathbb{Z}_{p}}\text{log}_{p}d(x,y)d\mu(y)$
$=$ $\int_{\mathbb{Z}_{p}}\text{log}_{p}(p^{-N}|t|_{p})p^{-N}dt$
$=$ $-\left(N+\frac{1}{p-1}\right)p^{-N}$
and
$\int_{B_{\epsilon}(x)}d(x,y)^{n}d\mu(y)=$ $\int_{x+p^{v_{p}(x)+N}\mathbb{Z}_{p}}d(x,y)^{n}d\mu(y)$
$=$ $\int_{\mathbb{Z}_{p}}(p^{-N}|t|_{p})^{n}p^{-N}dt$
$=$ $\frac{p^{n}(p-1)}{p^{n+1}-1}p^{-N(n+1)}.$
So we have
$\int_{B_{\epsilon}(x)}|G(x,y)|d\mu(y)\leq CNp^{-N}\leq C^{\prime}\epsilon|\ln\epsilon|,$
thus $G\in L^{1}(X_{m})$. Additionally, as $\rho e^{u_{d}(y)}+\frac{\rho}{V}$ is bounded and $|\widetilde{G_{d}}(\cdot,y)|\leq|G(\cdot,y)|$ [14], we have
[…]
so $v_{d}\in L^{\infty}(X_{m,d})$, thus $\widetilde{v_{d}}\in L^{\infty}(X_{m})$.
@@
$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}=$ $\int_{X_{m}}\widetilde{D_{d}u_{d}}\phi d\widetilde{\mu_{d}}$
$=$ $\int_{X_{m}}(D_{d}\widetilde{u_{d}})\phi d\widetilde{\mu_{d}}$
$=$ $\int_{X_{m}}(D_{d}\phi)\widetilde{u_{d}}d\widetilde{\mu_{d}}$
@@
[…]
$\leq$ $\rho\int_{B_{\epsilon}(y)}|D_{d}\phi(x)||\widetilde{G_{d}}(x,y)|d\widetilde{\mu_{d}}(x)+\int_{B_{\epsilon}(y)}|D_{d}\phi||\widetilde{v_{d}}|d\widetilde{\mu_{d}}$
$\leq$ $C_{1}\epsilon|\ln\epsilon|+C_{2}\epsilon$
@@
$\mathrm{II}=\int_{X_{m}\backslash B_{\epsilon}(y)}(D_{d}\phi)\widetilde{u_{d}}d\widetilde{\mu_{d}}=\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)\widetilde{u_{d}}d\widetilde{\mu_{d}}$
for sufficiently large $d$.
And in $X_{m}\backslash B_{\epsilon}(y)$, since $\widetilde{u_{d}}(x)$ depend only on $|x|_{p}$, $|y|_{p}$ and $|x-y|_{p}$, so $\forall\ x_{0}\in X_{m}\backslash B_{\epsilon}(y)$, $\exists\ r_{0}(x_{0})<\min\{|x_{0}|_{p},|x_{0}-y|_{p}\}$, s.t. $\forall\ x_{0}\in B(x_{0},r_{0}(x_{0}))$, we have $|x|_{p}=|x_{0}|_{p}$, $|x-y|_{p}=|x_{0}-y|_{p}$. So $\widetilde{u_{d}}(x)$ is locally constant on $B(x_{0},r_{0}(x_{0}))$. Since $X_{m}\backslash B_{\epsilon}(y)$ is compact, so $X_{m}\backslash B_{\epsilon}(y)=\bigcup_{i=1}^{K}B(x_{i},r_{i}(x_{i}))$. Taking $\delta=\min\{r_{1}(x_{1}),\dots,r_{K}(x_{K})\}$, then $\forall\ x,z\in X_{m}\backslash B_{\epsilon}(y)$ with $|x-z|_{p}<\delta$, we have $\widetilde{u_{d}}(x)=\widetilde{u_{d}}(z)$ as balls on $p$-adic field are either disjoint or concentric. And we notice that $|\widetilde{u_{d}}|\leq C_{\epsilon}$, so there exists a subsequence $\widetilde{u_{d_{k}}}$ converges to a locally constant function $u$ on $X_{m}\backslash B_{\epsilon}(y)$ such that
$\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)\widetilde{u_{d_{k}}}d\widetilde{\mu_{d_{k}}}=\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)(\widetilde{u_{d_{k}}}-u)d\widetilde{\mu_{d_{k}}}+\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)ud\widetilde{\mu_{d_{k}}}$.
The first term can be controlled by $\widetilde{u_{d_{k}}}-u||_{L^{\infty}(X_{m}\backslash B_{\epsilon}(y))}$, so
$\mathrm{II}\rightarrow\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)ud\mu.$
Letting $\epsilon\rightarrow 0$ and by the standard diagonal argument, we obtain a convergent function $u$, so
$\int_{X_{m}}\widetilde{D_{d_{k}}u_{d_{k}}}\phi d\widetilde{\mu_{d_{k}}}\rightarrow\int_{X_{m}}(D\phi)ud\mu=\int_{X_{m}}(Du)\phi d\mu$
@@
What’s more, since
$\int_{X_{m}}\phi d\nu=\lim_{k\to\infty}\int_{X_{m}}\phi e^{\widetilde{u_{d_{k}}}}d\widetilde{\mu_{d_{k}}}=\int_{X_{m}}\phi e^{u}d\mu,$
and by definition, we have
$\int_{X_{m}}\phi d\nu=\int_{X_{m}}\phi e^{\bar{u}}d\mu,$
so $\bar{u}=u$ for a.e. $x$. Then by (4.4) we get
$\int_{X_{m}}(Du)\phi d\mu+\rho\int_{X_{m}}e^{u}\phi d\mu=\rho\phi(y).$
@@
Since the solution we found converges from the solution on finite quotient, it inherits the properties of the solution on finite quotient. So the solution $u$ has those radially symmetric and monotone properties with a universal upper bound $M$.
When $\rho\geq\rho^{*}$, then there may exist non-monotone radially symmetric solutions, let alone a non-radially symmetric one, so that $u$ will not have the universal upper bound, this may prevent the convergence of solutions on the finite quotient, and hence the existence cannot be guaranteed.
@@ 4.2 Uniqueness of the solution
Although the solution on finite quotient is unique when $\rho\in(0,\rho^{*})$, we can’t ensure uniqueness on the Tate curve as there may has another solution on the Tate curve but vanishes on finite quotient. We now address uniqueness on the Tate curve by a standard method [20]. First, we show that the solution is unique when $\rho\leq\varepsilon$. Next, we prove the linearized solution is non-degenerate under the same constraints of $\rho$, then we can get the uniqueness.
@@
$D\phi_{n}+V_{n}(x)\phi_{n}=0,$
where $V_{n}(x)=\rho_{n}\frac{e^{u_{n}}-e^{v_{n}}}{u_{n}-v_{n}}$. Since $u(x)\leq M$ by Theorem 3.13, $V_{n}(x)$ converges to 0. And it is easy to see that $\phi_{n}$ uniformly converges to the eigenfunction of $D$ with the eigenvalue 0, which is the harmonic function on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$. As $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ is compact without boundary, then the harmonic function on it must be a constant function, which contradicts the fact that $\phi_{n}$ must change the sign. ∎
@@
$Df+\rho e^{u}f=0.$ (4.5)
If (4.5) is nondegenerate for $\rho\leq\rho^{*}$, then Lemma 4.2 implies the uniqueness of the solution for such $\rho$. To address it, we give the following Lemma:
@@
$\mu_{k}<0<\mu_{k+1},$
which means 0 is not the eigenvalue of LV(x)L-V(x). Thus the operator LV(x)L-V(x) is nondegenerate. ∎
Theorem 4.4.
When $\rho\in(0,\rho^{*})$, then the solution of (3.1) is unique.
Proof.
By Theorem 3.13, we can see that $0<\rho e^{u}<\rho^{*}e^{M}=\theta_{p,m}$. And from Theorem 2.4 and Lemma 3.15, the smallest nonzero eigenvalue of D-D is
$\min_{\begin{subarray}{c}1\leq k\leq m-1\\ i\geq 1\\ 0\leq j\leq m-1\end{subarray}}\left\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\right\}$ (4.6)
$=$ $\min\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
@@
$=$ $\theta_{p,m}.$
Then from Lemma 4.3, the linearized operator is nondegenerate, then we obtain the uniqueness. ∎